14. 15t^2-17t-4
(5t-3)(t-4)
Tried to work it out don't know if this is the right answer I do not get how to doe these. Can you please check thank you very much I don't know what I would do with out such good help like you guy and gals. Thanks again

Answer by oberobic(2304)

You're close. Try: (5t + 1)(3t - 4).
Using the distributive property, you will obtain:
15t^2 - 20t + 3t - 4, which simplifies to 15t^2 - 17t - 4.
Then solve for t by setting each factor to zero. That is,
(5t + 1) = 0 and (36t - 4) = 0.
For 5t + 1 to equal zero, then 5t must = -1 (i.e., -1 + 1 = 0), so t = -1/5.
For 3t - 4 to equal zero, then 3t must equal 4 (i.e., 4 - 4 = 0), so t = 4/3.
Therefore, your solution set might be t = { -1/5, 4/3 }
I say "might be" because there may be a constraint.
If t cannot be a negative number, then your only solution is 4/3.
Likewise, if t cannot be a positive number, then your only solution is -1/5.